OpenStudy (anonymous):

An airplane has a total of 130 seats. The number of coach-class seats is 20 more than two times the number of first-class seats. How many of each type of seat are there on the plane? please show work

OpenStudy (anonymous):

could you do a check step please

OpenStudy (anonymous):

The question is wrong. Check the question.

OpenStudy (shane_b):

I reviewed what I posted earlier and saw a silly mistake (I've done a lot of these today..sorry). Here it is redone: Let C = coach seats and F = first class seats: \[C-2F=20\]\[C+F=130\]Solve for one of the unknowns in terms of the other: \[F=130−C\]Plug that back into the first equation: \[C-2(130-C)=20\]Solve for C to get 93.33 seats. Now that you know C you can solve for F using one of the equations above: \[2F+20=93.33\]\[F=36.67\] As you can see...something is wrong with the original question since the number of coach and first class seats came out to be fractional values.